# TREAD: Token Routing for Efficient Architecture-agnostic Diffusion Training

### v2

@@ Abstract
Diffusion models have emerged as the mainstream approach for visual generation. However, these models typically suffer from sample inefficiency and high training costs. Consequently, methods for efficient finetuning, inference and personalization were quickly adopted by the community. However, training these models in the first place remains very costly. While several recent approaches — including masking, distillation, and architectural modifications — have been proposed to improve training efficiency, each of these methods comes with a tradeoff: they achieve enhanced performance at the expense of increased computational cost or vice versa. In contrast, this work aims to improve training efficiency as well as generative performance at the same time through routes that act as a transport mechanism for randomly selected tokens from early layers to deeper layers of the model. Our method is not limited to the common transformer-based model — it can also be applied to state-space models and achieves this without architectural modifications or additional parameters. Finally, we show that TREAD reduces computational cost and simultaneously boosts model performance on the standard ImageNet-256 benchmark in class-conditional synthesis. Both of these benefits multiply to a convergence speedup of 14 at 400K training iterations compared to DiT and 37 compared to the best benchmark performance of DiT at 7M training iterations. Furthermore, we achieve a competitive FID of 2.09 in a guided and 3.93 in an unguided setting, which improves upon the DiT, without architectural changes. Find our code at: https://github.com/CompVis/tread.
@@ 1 Introduction
In recent years, diffusion models [sohl2015deep, ho2020denoising, rombach2022high_latentdiffusion_ldm] have become a powerful generative technique for image synthesis [rombach2022high_latentdiffusion_ldm]. They have also been successfully extended to the 3D [poole2023dreamfusion] and video domains [blattmann2023align_videoldm]. While diffusion models avoid some training challenges faced by their predecessors, such as GANs [goodfellow2020generative], which are inherently unstable during training, they incur high costs [liang2024scalinglawsdiffusiontransformers] due to slow convergence and sample inefficiency [mei2023deepnetworksdenoisingalgorithms]. Currently, the Diffusion Transformer (DiT) [dit_peebles2022scalable] is the main backbone when training diffusion models, building on the established transformer architecture [vaswani2017attention]. However, the Transformer architecture has computational cost limitations, as it scales quadratically with token length and converges slowly, further increasing the cost of training diffusion models. Training a DiT on standard benchmarks alone requires thousands of A100 GPU hours without reaching convergence, and text-to-image models demand even more resources — Stable Diffusion [rombach2022high_latentdiffusion_ldm], for instance, required about 150,000 A100 GPU hours. Despite ongoing efforts to reduce computational demands through improved infrastructure, implementation, and hardware, democratizing the training of diffusion models remains a distant goal. Methods such as LoRA [hu2022lora] for parameter-efficient fine-tuning, caching strategies [wimbauer2024cache, ma2024deepcache] to accelerate inference, and approaches for personalization [gal2022image, hulfm, ruiz2023dreambooth, baumann2024continuous] have made these aspects more accessible. However, full-scale training remains extremely costly, limiting broad participation in the development and improvement of diffusion models. Several works have proposed methods to accelerate training convergence utilizing external self-supervised features [hu2023self, yu2024repa], improved flow-based theory [lipman2022flow, gui2024depthfm, schusterbauer2024boosting], or optimized data combinations [tong2023improving, mixup]. Another approach to reducing computational resource requirements is masking, where a subset of the available information is used during training. This subset can be selected either randomly [zheng2023fast_maskdit, Gao_2023_ICCV] or based on learned heuristics [wang2024attentiondriventrainingfreeefficiencyenhancement, meng2022adavit], while the remaining information is permanently lost [devlin2018bert, he2022masked_mae]. Beyond masking, computational efficiency can also be improved by controlling the flow of information during training. Instead of discarding information, it can be selectively directed to specific computational blocks, allowing tokens to skip computations and reducing the overall workload. This process is known as routing [raposo2024mixture].
@@
A novel training strategy for token-based sequential models that utilizes a token transport mechanism termed route, which directs tokens from early into later network layer.
@@ 2 Related Work
Score-based models [song2019generative, song2020improved], such as Denoising Diffusion Probabilistic Models (DDPM) [ho2020denoising], have become the dominant framework for image synthesis in recent years. These models start with a Stochastic Differential Equation (SDE) and gradually add Gaussian noise to real data in the forward process. This process transforms the real distribution into a Gaussian distribution. An iterative denoising process is then used to reverse this transformation and generate samples. A key improvement in the efficiency of early diffusion models was to train in a compressed latent space [rombach2022high_latentdiffusion_ldm].
Building on the foundation of score-based models, early diffusion models [ho2020denoising, dhariwal2021diffusion, rombach2022high_latentdiffusion_ldm] used the UNet architecture [unet] as their backbone. However, more recently, token-based models such as DiT [dit_peebles2022scalable] have become a preferred backbone choice [zheng2023fast_maskdit, esser2024scalingrectifiedflowtransformers, chen2023pixartalphafasttrainingdiffusion]. However, these models face challenges due to quadratic complexity w.r.t. the number of tokens.
To address the computational challenges associated with token-based models, recent studies have proposed caching mechanisms to accelerate diffusion models based on UNets [ma2023deepcache] or DiTs [ma2024learningtocacheacceleratingdiffusiontransformer]. Unlike these approaches, our routing scheme does not skip the computation of network layers for each token. Instead, we introduce a routing mechanism that is applied only during training. This mechanism transports tokens from one network layer to another, enhancing the model’s efficiency without omitting necessary computations.
In addition to diffusion transformers, State Space Models (SSMs) have recently been shown to be promising alternatives to DiTs [hu2024zigma, fei2024scalablediffusionmodelsstate, fei2024diffusion, teng2024dimdiffusionmambaefficient] to alleviate the quadratic complexity. However, largely because of their recent rise, there are currently few works that push the efficiency of SSMs in a generative setting. Token pruning [zhan2024exploringtokenpruningvision], token masking [tang2024mambamimpretrainingmambastate], and mixture-of-experts [pioro2024moemambaefficientselectivestate] for SSMs have been explored so far, although not for diffusion backbones. Unlike them, we consider the token routing technique in state-space models to improve training efficiency.
@@
Training diffusion models is challenging due to their sample inefficiency and high computational demands. Two primary factors must be balanced during optimization: training cost (iterations per second) and training effectiveness (performance gain per iteration). Increasing iterations per second reduces the cost per iteration, while improved training effectiveness lowers the overall compute time by achieving higher quality outputs in fewer steps. Many methods optimize one factor at the expense of the other. For example, masking a subset of tokens [zheng2023fast_maskdit] significantly enhances throughput in transformer networks, which are hindered by quadratic complexity , and similar benefits are observed in SSMs [li2024videomamba]. Alternatively, training effectiveness is sometimes enhanced via architectural changes (e.g., additional parameters [zheng2023fast_maskdit, tang2024generative]) or pretrained teacher models for representation distillation [yu2024repa]. However, these architecture modifications generally compromise training efficiency as more parameters or additional models are required during training, whereas masking-based methods that enhance training effectiveness [devlin2018bert, he2022masked_mae, Gao_2023_ICCV] are limited by the inherent drawback of permanent data loss.
@@
Residual-based architectures, including transformers [resnet], can interpret outputs from preceding layers [raposo2024mixture, ma2023deepcache, he2024matters, ma2024learningtocacheacceleratingdiffusiontransformer]. We demonstrate this property in Figure 4, where the cosine similarity between the outputs of all layers in a trained network is shown. This characteristic can be leveraged either to cache layer outputs from previous timesteps for improved inference or to route tokens from one layer to deeper layers, as proposed in TREAD. While diffusion models are typically trained using the same sequence of layers for all samples for each denoising step, TREAD introduces additional variance. Specifically, the second pathway via route reduces the number of transformations that are applied to all tokens by the route length. As tokens are randomly selected for routing, the network learns the denoising step with a variable number of layers. Since the network always has access to the standard full pathway (the selection rate for is less than 1.0), it is beneficial to design the route such that the alternative pathway differs clearly from the standard one. The extent of this difference depends on the route length. Longer routes produce more pronounced variations. This is reflected in the training loss as longer routes correlate with increased training loss compared to the baseline without routing. Good generative performance and low FID (see Figure 6) also correlate empirically strongly with route length and, therefore also with the increase in loss. Consequently, longer routes improve both training efficiency and training effectiveness, yielding multiplicative speed-ups (iterations per second effectiveness per iteration).
@@ 4.1 Experimental Details
Our model follows the standard two-stage training process of Latent Diffusion Models (LDM) [rombach2022high_latentdiffusion_ldm], where an autoencoder is first trained to map between pixel-space and a compressed latent representation. We utilize the pre-trained VAE from Stable Diffusion with a standard downsampling factor of 8. As our method primarily requires token-based processing, we adopt a standard DiT-XL architecture [dit_peebles2022scalable] with a patch size of 2 as our main model. Additionally, we demonstrate the generalizability of our approach to other architectures, such as Diffusion-RWKV [fei2024diffusion]. For ablations we use a DiT-B/2 if not otherwise specified. All modifications are explicitly stated for each experiment. To indicate the application of our method, we append +TREAD to the respective models.
@@
We follow the standard of existing works on diffusion models and train models on ImageNet-256 with a batch size of . Furthermore, we follow the standard setting for AdamW from the original DiT [dit_peebles2022scalable] with a learning rate of 1e-4 and no weight decay and the standard -parameters of and respectively. We use the standard EMA approach with an update rate of . All results are computed using this EMA model. We do not employ any kind of data augmentation. All experiments are conducted on GB-A100 GPUs with a local batch size of 64, resulting in the standard global batch size of 256. In addition, we use precomputed latents of ImageNet-256, leading to a representation. Except for experiments that specifically state otherwise, we apply a selection rate of to our routing mechanism. For Transformer-based structures, we use random selection, while for sequential models like RWKV, we find that row selection works better.
@@ 4.2 Main Result
We demonstrate the effectiveness of our approach in improving efficiency and performance while showcasing its adaptability to diverse architectures. To illustrate this, we use RWKV [peng2023rwkv], which shares many aspects of SSMs and employs a RNN-like structure, which makes it fundamentally different from the transformer-based DiT. As shown in Table 2, our method consistently outperforms the baseline models across both architectures. Notably, the performance gains are also pronounced in larger models, indicating the scalability of our approach and its suitability for larger architectures. We also expand TREAD beyond EDM [karras2022elucidating] to flowmatching [lipman2022flow, ma2024sitexploringflowdiffusionbased, albergo2022building] and find that our method transfers well to other frameworks. All models, except those denoted as ”SiT” in Table 2 are trained using EDM to remain comparable to our baselines and competitors. We also extend our experiments beyond ImageNet-256 to ImageNet-512 to investigate the impact of larger token counts (see Table 4).
Model #Param Iter. FID DiT-S/2 [dit_peebles2022scalable] 33M 400K 68.40 + TREAD 47.04▾21.36 SiT-S/2 [ma2024sit] 57.60 + TREAD 38.08▾19.52 DiT-B/2 [dit_peebles2022scalable] 130M 400K 43.47 + TREAD 21.34▾22.13 SiT-B/2 [ma2024sit] 33.00 + TREAD 15.98▾16.02 DiT-L/2 [dit_peebles2022scalable] 458M 400K 23.33 + TREAD 9.10▾14.23 SiT-L/2 [ma2024sit] 18.80 + TREAD 5.91▾12.89 DiT-XL/2 [dit_peebles2022scalable] 675M 400K 19.47 + TREAD 7.38▾12.09 SiT-XL/2 [ma2024sit] 17.20 + TREAD 4.89▾12.31 RWKV [fei2024diffusion] 210M 400K 59.84 + TREAD 53.79▾6.05
@@
Scaling to Larger Resolutions and Text-to-Image Datasets.
We also investigate the scalability and general applicability of TREAD by validating it in higher-resolution and more general settings. Specifically, we evaluate on ImageNet-512 and MS-COCO [coco]. We train DiT-B/2 models with and without TREAD for 400K iterations each and evaluate using FID@10k for both. We find that our method leads to substantial gains in generative performance, reducing FID by more than a third, while also leading to substantially faster training due to a significant increase in iteration speed, as indicated in Table 4.
Model #Param Iter. FID@10K ImageNet-512 DiT 62.8 + TREAD 130M 400K 40.9▾21.9 MS-COCO DiT 35.68 + TREAD 130M 400K 20.55▾15.13
@@ 4.3 Ablation Study
In Figure 8, we present results from a set of DiT-B/2 models trained with different routes . We conduct two ablations: first, by fixing and varying the end layer , and second, by fixing and varying the starting layer . From these experiments, we observe two distinct modes: one where routing improves performance and one where it does not. Since these modes are clearly separable, we can derive principles for route placement to enhance training effectiveness. Specifically, we find that i) longer routes generally lead to better performance, ii) starting layers up to improve results, and iii) ending layers up to are beneficial. Additionally, the configuration demonstrates that choosing an excessively late end location can be detrimental to performance, as the network requires sufficient capacity to integrate routed tokens into the overall information flow. Whenever routing does not lead to lower FID, we can attribute the observed effect to violating at least one of these principles. In the case of TREAD, we can optimize both training effectiveness and training efficiency simultaneously. According to our route placement guidelines, training effectiveness can be increased with longer routes. At the same time, longer routes accelerate training speed by increasing the number of layers processing fewer tokens compared to DiT [dit_peebles2022scalable]. In Figure 6, we present our findings from this analysis. Our results indicate that route length is strongly correlated with both FID and training loss, even at early stages. Consequently, early training loss also exhibits a strong correlation with FID. However, unlike standard diffusion training, a lower loss does not necessarily indicate better performance. Instead, we observe the opposite: a higher difference in loss with respect to the baseline reflects the greater influence of the routing mechanism as discussed in Section 3, which in turn relates to improved performance. This can help gauge the routing mechanism’s impact on the diffusion model at early training stages.
Furthermore, we also demonstrate the importance of the start layer and end layer in Figure 7, where we show the generative performance for models with the respective settings. Notably, there is a trend for both where the best results were achieved using and , with deviations performing similarly in FID (except for ).
@@
Selection Rate FID 0 (DiT Baseline) 43.47 0.25 26.50▾16.97 0.50 21.34▾22.13 0.75 49.29▴5.82
Model FID DiT-XL/2 +TREAD (relative) 18.81 +TREAD (absolute) 7.38

### v3

@@ Abstract
Diffusion models have emerged as the mainstream approach for visual generation. However, these models typically suffer from sample inefficiency and high training costs. Consequently, methods for efficient finetuning, inference and personalization were quickly adopted by the community. However, training these models in the first place remains very costly. While several recent approaches including masking, distillation, and architectural modifications have been proposed to improve training efficiency, each of these methods comes with a tradeoff: they achieve enhanced performance at the expense of increased computational cost or vice versa. In contrast, this work aims to improve training efficiency as well as generative performance at the same time through routes that act as a transport mechanism for randomly selected tokens from early layers to deeper layers of the model. Our method is not limited to the common transformer-based model - it can also be applied to state-space models and achieves this without architectural modifications or additional parameters. Finally, we show that TREAD reduces computational cost and simultaneously boosts model performance on the standard ImageNet-256 benchmark in class-conditional synthesis. Both of these benefits multiply to a convergence speedup of 14 at 400K training iterations compared to DiT and 37 compared to the best benchmark performance of DiT at 7M training iterations. Furthermore, we achieve a competitive FID of 2.09 in a guided and 3.93 in an unguided setting, which improves upon the DiT, without architectural changes.
Project Page: https://compvis.github.io/tread
@@ 1 Introduction
In recent years, diffusion models [sohl2015deep, ho2020denoising, rombach2022high_latentdiffusion_ldm] have become a powerful generative technique for image synthesis [rombach2022high_latentdiffusion_ldm]. They have also been successfully extended to the 3D [poole2023dreamfusion] and video domains [blattmann2023align_videoldm]. While diffusion models avoid some training challenges faced by their predecessors, such as GANs [goodfellow2020generative], which are inherently unstable during training, they incur high costs [liang2024scalinglawsdiffusiontransformers] due to slow convergence and sample inefficiency [mei2023deepnetworksdenoisingalgorithms]. Currently, the Diffusion Transformer (DiT) [dit_peebles2022scalable] is the main backbone when training diffusion models, building on the established Transformer architecture [vaswani2017attention]. However, the Transformer architecture has computational cost limitations, as it scales quadratically with token length and converges slowly, further increasing the cost of training diffusion models. Training a DiT on standard benchmarks alone requires thousands of A100 GPU hours without reaching convergence, and text-to-image models demand even more resources. Stable Diffusion [rombach2022high_latentdiffusion_ldm], for instance, required about 150,000 A100 GPU hours. Despite ongoing efforts to reduce computational demands through improved infrastructure, implementation, and hardware, democratizing the training of diffusion models remains a distant goal. Methods such as LoRA [hu2022lora] for parameter-efficient fine-tuning, caching strategies [wimbauer2024cache, ma2024deepcache] to accelerate inference, and approaches for personalization [gal2022image, hulfm, ruiz2023dreambooth, baumann2024continuous] have made these aspects more accessible. However, full-scale training remains extremely costly, limiting broad participation in the development and improvement of diffusion models. Several works have proposed methods to accelerate training convergence utilizing external self-supervised features [hu2023self, yu2024repa, fuest2024diffusion], improved flow-based theory [lipman2022flow, gui2024depthfm, schusterbauer2024boosting], or optimized data combinations [tong2023improving, mixup]. Another approach to reducing computational resource requirements is masking, where a subset of the available information is used during training. This subset can be selected either randomly [zheng2023fast_maskdit, Gao_2023_ICCV] or based on learned heuristics [wang2024attentiondriventrainingfreeefficiencyenhancement, meng2022adavit], while the remaining information is permanently lost [devlin2018bert, he2022masked_mae]. Beyond masking, computational efficiency can also be improved by controlling the flow of information during training. Instead of discarding information, it can be selectively directed to specific computational blocks, allowing tokens to skip computations and reducing the overall workload. This process is known as routing [raposo2024mixture].
@@
A novel training strategy for token-based sequential models that utilizes a token transport mechanism termed routing, which directs tokens from early into later network layer.
@@ 2 Related Work
Score-based generative models [song2019generative, song2020improved], like notably DDPMs [ho2020denoising], now lead image synthesis. They define a forward SDE that progressively adds Gaussian noise, mapping data to a standard Gaussian. An iterative denoising process is used to reverse this transformation and generate samples. A key improvement in the efficiency of early diffusion models was to train in a compressed latent space [rombach2022high_latentdiffusion_ldm]. Building on the foundation of score‑based models, early diffusion methods [ho2020denoising, dhariwal2021diffusion, rombach2022high_latentdiffusion_ldm] used the UNet backbone [unet], while more recent token‑based architectures like DiT [dit_peebles2022scalable] and its variants [zheng2023fast_maskdit, esser2024scalingrectifiedflowtransformers, chen2023pixartalphafasttrainingdiffusion] have gained preference, despite their quadratic complexity in the number of tokens. To address computational challenges associated with token-based models, recent work introduces caching for UNet‑based diffusion [ma2023deepcache] or DiT‑based models [ma2024learningtocacheacceleratingdiffusiontransformer]. Unlike these approaches, our method instead applies a training‑time routing mechanism that moves tokens between layers, improving efficiency without omitting necessary computations. Alongside diffusion transformers, state‑space models (SSMs) have emerged as promising alternatives to DiTs [hu2024zigma, fei2024scalablediffusionmodelsstate, fei2024diffusion, teng2024dimdiffusionmambaefficient] to mitigate quadratic complexity. Explorations of token pruning [zhan2024exploringtokenpruningvision], token masking [tang2024mambamimpretrainingmambastate] and mixture‑of‑experts [pioro2024moemambaefficientselectivestate] for SSMs have not targeted diffusion backbones, whereas we consider token routing in SSMs aswell to boost overall training efficiency and convergence.
@@
TREAD: Our method for efficient diffusion training.
In a) the standard training and inference strategy is shown where all tokens are processed by all layers of the network. TREAD enhances training efficiency by routing tokens around certain layers by reducing computational load and it preserves information which is shown in b). Since TREAD is used only during training, the standard setting shown in a) is used for inference.
@@
Balancing Cost and Effectiveness.
Training diffusion models is challenging due to their sample inefficiency and high computational demands. Two primary factors must be balanced during optimization: training cost (iterations per second) and training effectiveness (performance gain per iteration). Increasing iterations per second reduces the cost per iteration, while improved training effectiveness lowers the overall compute time by achieving higher quality outputs in fewer steps. Many methods optimize one factor at the expense of the other, creating a trade-off between cost and effectiveness. For example, masking a subset of tokens [zheng2023fast_maskdit] significantly enhances throughput in transformer networks, which are hindered by quadratic complexity , and similar benefits are observed in SSMs [li2024videomamba]. Alternatively, training effectiveness is sometimes enhanced via architectural changes (e.g., additional parameters [zheng2023fast_maskdit, tang2024generative]) or pretrained teacher models for representation distillation [yu2024repa]. However, these architecture modifications generally compromise training efficiency as more parameters or additional models are required during training, whereas masking-based methods that enhance training effectiveness [devlin2018bert, he2022masked_mae, Gao_2023_ICCV] are limited by the inherent drawback of permanent data loss.
@@
Residual-based architectures, including transformers [resnet], can interpret outputs from preceding layers [raposo2024mixture, ma2023deepcache, he2024matters, ma2024learningtocacheacceleratingdiffusiontransformer]. We demonstrate this property in Figure 4, where the cosine similarity between the outputs of all layers in a trained network is shown. This characteristic can be leveraged either to cache layer outputs from previous timesteps for improved inference or to route tokens from one layer to deeper layers, as proposed in TREAD. While diffusion models are typically trained using the same sequence of layers for all samples for each denoising step, TREAD introduces additional variance. Specifically, the second pathway via route reduces the number of transformations that are applied to all tokens by the route length. As tokens are randomly selected for routing, the network learns the denoising step with a variable number of layers. This implicitly motivates the network to produce less washed-out features in layer as they are later being used by layer directly. Since the network always has access to the standard full pathway (the selection rate for is less than 1.0), it is beneficial to design the route such that the alternative pathway differs clearly from the standard one as this will increase the above mentioned effect. Good generative performance and low FID (see Figure 6) also correlate strongly with route length. Consequently, longer routes improve both training efficiency and training effectiveness, yielding multiplicative speed-ups (iterations per second effectiveness per iteration).
@@ 4.1 Experimental Details
Our model follows the standard two-stage training process of Latent Diffusion Models (LDM) [rombach2022high_latentdiffusion_ldm], where an autoencoder is first trained to map between pixel-space and a compressed latent representation. We utilize the pre-trained VAE from Stable Diffusion with a standard downsampling factor of 8. As our method primarily requires token-based processing, we adopt a standard DiT-XL architecture [dit_peebles2022scalable] with a patch size of 2 as our main model. Additionally, we demonstrate the generalizability of our approach to other architectures, such as Diffusion-RWKV [fei2024diffusion] and Mamba [gu2023mamba]. For ablations we use a DiT-B/2 if not otherwise specified. All modifications are explicitly stated for each experiment. To indicate the application of our method, we append +TREAD to the respective models.
@@
We follow the standard of existing works on diffusion models and train models on ImageNet-256 with a batch size of . Furthermore, we follow the standard setting for AdamW from the original DiT [dit_peebles2022scalable] with a learning rate of 1e-4 and no weight decay and the standard -parameters of and respectively. We use the standard EMA approach with an update rate of . All results are computed using this EMA model. We do not employ any kind of data augmentation. All experiments are conducted on GB-A100 GPUs with a local batch size of 64, resulting in a global batch size of 256, which reflects the standard for ImageNet-256. In addition, we use precomputed latent representations of ImageNet-256, leading to a representation. Except for experiments that specifically state otherwise, we apply a selection rate of to our routing mechanism. For Transformer-based structures, we use random selection, while for sequential models like RWKV and Mamba, we find that row selection works better.
@@ 4.2 Main Result
We demonstrate the effectiveness of our approach in improving efficiency and performance while showcasing its adaptability to diverse architectures. To illustrate this, we use RWKV [peng2023rwkv], which shares many aspects of SSMs and employs a RNN-like structure, which makes it fundamentally different from the transformer-based DiT. As shown in Table 2, our method consistently outperforms the baseline models across both architectures. Notably, the performance gains are also pronounced in larger models, indicating the scalability of our approach and its suitability for larger architectures. We also expand TREAD beyond EDM [karras2022elucidating] to flowmatching [lipman2022flow, ma2024sitexploringflowdiffusionbased, albergo2022building] and find that our method transfers well to other frameworks. All models, except those denoted as “SiT” in Table 2 and Table 5 are trained using EDM to remain comparable to our baselines and competitors. We also extend our experiments beyond ImageNet-256 to ImageNet-512 to investigate the impact of larger token counts (see Table 4).
Model #Param Iter. Batchsize FID DiT-S/2 [dit_peebles2022scalable] 33M 400K 256 68.40 + TREAD 33M 400K 256 47.04▾21.36 SiT-S/2 [ma2024sit] 33M 400K 256 57.60 + TREAD 33M 400K 256 38.08▾19.52 DiT-B/2 [dit_peebles2022scalable] 130M 400K 256 43.47 + TREAD 130M 400K 256 21.34▾22.13 SiT-B/2 [ma2024sit] 130M 400K 256 33.00 + TREAD 130M 400K 256 15.98▾16.02 DiT-L/2 [dit_peebles2022scalable] 458M 400K 256 23.33 + TREAD 458M 400K 256 9.10▾14.23 SiT-L/2 [ma2024sit] 458M 400K 256 18.80 + TREAD 458M 400K 256 5.91▾12.89 DiT-XL/2 [dit_peebles2022scalable] 675M 400K 256 19.47 + TREAD 675M 400K 256 7.38▾12.09 SiT-XL/2 [ma2024sit] 675M 400K 256 17.20 + TREAD 675M 400K 256 4.89▾12.31 LightningDiT [yao2025reconstruction] 675M 80K 1024 5.14 + TREAD 675M 80K 1024 3.71▾1.43 RWKV [fei2024diffusion] 210M 400K 256 59.84 + TREAD 210M 400K 256 53.79▾6.05 Mamba [maml] 111M 400K 256 69.39 + TREAD 111M 400K 256 61.17▾8.22
@@
Scaling to Larger Resolutions and Text-to-Image.
We also investigate the scalability and general applicability of TREAD by validating it in higher-resolution and more general settings. Specifically, we evaluate on ImageNet-512 and T2I MS-COCO (2014) [coco]. We train DiT-B/2 models with and without TREAD for 400K iterations each using a batchsize of 256 and evaluate using FID@10k for both. We find that our method leads to substantial gains in generative performance, reducing FID by more than a third (as indicated in Table 4), while also leading to substantially faster training due to a significant increase in iteration speed.
Model #Param Iter. Batchsize FID@10K ImageNet-512 DiT 130M 400K 256 62.88 + TREAD 130M 400K 256 40.91▾21.97 MS-COCO DiT 130M 400K 256 35.68 + TREAD 130M 400K 256 20.55▾15.13
@@ 4.3 Ablation Study
In Figure 8, we present results from a set of DiT-B/2 models trained with different routes . We conduct two ablations: first, by fixing and varying the end layer , and second, by fixing and varying the starting layer . From these experiments, we observe two distinct modes: one where routing improves performance and one where it does not. Since these modes are clearly separable, we can derive principles for route placement to enhance training effectiveness. Specifically, we find that i) longer routes generally lead to better performance, ii) starting layers up to improve results, and iii) ending layers up to are beneficial. Additionally, the configuration demonstrates that choosing an excessively late end location can be detrimental to performance, as the network requires sufficient capacity to integrate routed tokens into the overall information flow. Whenever routing does not lead to lower FID, we can attribute the observed effect to violating at least one of these principles. In the case of TREAD, we can optimize both training effectiveness and training efficiency simultaneously. According to our route placement guidelines, training effectiveness can be increased with longer routes. At the same time, longer routes accelerate training speed by increasing the number of layers processing fewer tokens compared to DiT [dit_peebles2022scalable]. In Figure 6, we present our findings from this analysis. Our results indicate that route length is strongly correlated with both FID and training loss, even at early stages. Consequently, early training loss also exhibits a strong correlation with FID. Additionally to the usual diffusion loss we see an offset caused by the route that indicates its impact on the model. Furthermore, we provide a custom baseline in Table 5 where we compare TREAD against token-level dropout in each layer to underline the importance of route configuration.
We also demonstrate the importance of the start layer and end layer in Figure 7, where we show the generative performance for models with the respective route placement settings. Notably, there is a trend for both where the best results were achieved using and , with deviations following a clear downwards direction in terms of FID (except for ).
@@
| Selection Rate | FID |
| 0 (DiT-B/2) | 43.47 |
| 0.25 | 26.50▾16.97 |
| 0.50 | 21.34▾22.13 |
| 0.75 | 49.29▴5.82 |
| Model | FID |
| SiT-B/2 [ma2024sit] | 33.00 |
| TLDrop | 32.00▾1.00 |
| TLDropR | 24.28▾8.72 |
| +TREAD (absolute) | 15.98▾17.02 |
Model FID DiT-XL/2 19.47 +TREAD (relative) 18.81▾0.66 +TREAD (absolute) 7.38▾12.09